OpenStudy (anonymous):

A plane travels 350 km south along a straight path with an average velocity of 125 km/h to the south. The plane has a 30 minute layover, and then travels 220 km south with an average velocity of 115 km/h to the south. What is the average velocity for the trip?

OpenStudy (jack1):

v = distance / time so leg 1 of journey: V1 = Dist 1 / time 1 125 = 350 / time 1 2.8 hrs = time 1 leg 2: layover: time 2 = 30 mins = 0.5 hrs leg 3 : Veloc 3 = Dist 3 / Time 3 115 = 220 / time 3 time 3 = 1.91 hours

OpenStudy (jack1):

therefore: average velocity for whole trip = distance traveled for whole trip / time taken for whole trip V av = (D1 + D2 + D3) / (T1 + T2 + T3) V av = (350 + 0 + 220) / (2.8 + 0.5 + 1.9) V av = 570 km / 5.21 hrs V av = ?

OpenStudy (jack1):

can you work it out from here?